OpenStudy (taemin):

Will give medal solve the system by substitution (Show Work) -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

hartnn (hartnn):

3 equations. 3 unknowns what have you tried?

OpenStudy (taemin):

Input: {-x-y-z = -8, -4 x+4 y+5 z = 7, 2 x+2 z = 4}

OpenStudy (taemin):

solution x=3,y=6, z=-1 @hartnn I have the answer I need it to be checked

OpenStudy (taemin):

@AravindG can you help

OpenStudy (faiqraees):

eliminate y from the first two equations using elimination method

OpenStudy (photon336):

Pick a variable to eliminate it and Eliminate it in two equations. Say the first two equations. Then eliminate the same variable in the other two equations Say equations 1 and 3 Then you'll have two new equations you can use to find two of your variables.

OpenStudy (faiqraees):

then use anyone of the equation and then simultaneously solve it with the third one

OpenStudy (aravindg):

Just a tip: Try finding one variable in terms of other two from one of these equations. Substitute this variable in the next two equations. You get two equations with 2 unknowns. Now it can be solved easily.

OpenStudy (taemin):

ok thank you